OpenStudy (destinyyyy):

Help with 5 problems...

OpenStudy (destinyyyy):

2 [7- 3/2x] +5=21 [] are absolute value lines

OpenStudy (destinyyyy):

7- 3/2x =8

OpenStudy (destinyyyy):

Not sure what to do next

OpenStudy (destinyyyy):

@Nnesha @freckles @Vocaloid

Vocaloid (vocaloid):

you're on the right track, but when we have absolute value signs, we must consider the positive and negative solutions to the equation so, going by what you have so far, we have: | 7 - 3/2x | = 8, where | | are the absolute value signs, with me so far?

OpenStudy (destinyyyy):

Yes

Vocaloid (vocaloid):

So, going from there, we need to write two different equations 7 - 3/2x = 8 and 7 - 3/2x = -8

Vocaloid (vocaloid):

then we solve each one separately so, let's tackle the first one 7 - 3/2x = 8 first, subtract 7 from both sides and tell me what you get

OpenStudy (destinyyyy):

- 3/2x= 1

Vocaloid (vocaloid):

yup, now we divide both sides by (-3/2)

OpenStudy (destinyyyy):

um -2/3

Vocaloid (vocaloid):

yup, so x = -2/3 now, we're only halfway done, there's another value of x that works for the equation. now we solve 7 - (3/2)x = -8 using the same process as before. start by subtracting 7 from both sides

OpenStudy (destinyyyy):

x=10

Vocaloid (vocaloid):

yup, so x = 10 and x = -2/3 are your two solutions, good work ready for the next one?

OpenStudy (destinyyyy):

Yes thank you

OpenStudy (destinyyyy):

Solve. [x+8] +9=7 same as the one before

OpenStudy (destinyyyy):

I got x=-10

Vocaloid (vocaloid):

ah, not quite, there's an important rule to remember when it comes to absolute values. the absolute value of a mathematical expression cannot be negative as an example, if we have |x| = a, then a must be a positive number in order to have a solution so, looking at our problem, we first subtract 9 from each side to get |x + 8| = -2 the absolute value of any quantity must be positive, so this has no solution

OpenStudy (destinyyyy):

Okay

Vocaloid (vocaloid):

does that make sense? are you ready to move to the next one?

OpenStudy (destinyyyy):

Yes it does. I thought it was but wanted to make sure.. Solve using appropriate substitution. (x^2 -6x)^2 -11(x^2 -6x) -30=0

OpenStudy (destinyyyy):

I got 3+- square root 14 and 3+- square root 15 But its wrong...

Vocaloid (vocaloid):

huh, this is a tricky one, hold on a moment

OpenStudy (destinyyyy):

Okay

Vocaloid (vocaloid):

well, I'm not sure if this is the right approach but...I'm guessing it wants us to substitute a variable for x^2 -6x so, if we let a = x^2 -6x, then our equation becomes a^2 - 11a - 30 after we apply the quadratic formula we get